OpenStudy (anonymous):

Can someone help me on this one? 3x+6-x=8+3x-6. I think I need to combine the like terms first right? So 3x+3x+6+8-1x-6? Is this ok so far? Thank in advance for any help:)

OpenStudy (anonymous):

\[3x+6-x=8+3x-6\] \[2x+6=2+3x\] is a start

OpenStudy (anonymous):

2x + 6 = 2 + 3x --> 1x + 2 = 6 ---> x = 4

OpenStudy (anonymous):

Ok, I see, thanks so much:)